6 A truck with a speed of 20m/s slows down at a rate of 2.3m/2, what distance does it cover before stopping?
Answer #1

Initial velvety ce=comis Final velocidy ,v=omis accelerahon, a = - 2-3 mise v² = uz it has sa v?_uz 20 s = 0²- 202 24(+213) =